OpenStudy (anonymous):

Write a point slope equation for the line with the given slope that contains the given point. m=-3; (-3,-5)

OpenStudy (nenadmatematika):

y=-3x-14

OpenStudy (anonymous):

point-slope equation: \[y-y1=m(x-x1)\] \[y+5=-3(x+3)\]

OpenStudy (anonymous):

Thank you for the formulas!

OpenStudy (anonymous):

No problem